The Apiary still on EFB lockdown.   But the colonies still need a go through, and the grass needs cutting again.    Two custodians ,assisted by David Parkinson went through the colonies to check for Queen cells and any signs of further problems with EFB.  Temperature 18 degrees, Cloudy.
Hive# 1  Not gone through as 2 Qc’s left on f7 on 12/6,  no need to check till 3/7.
Hive#2  Inspected by David P.   Brood Chamber 14 x 12.   QS,  8 Qc’s taken down as unable to make up Nuc with 14×12 frames.  9 f brood BAS and eggs, stores ok.   Temper 8.   Super added above brood box as very little space for Q to lay up.
Hive#4  Inspected by Liz,    Brood chamber : National Deep.  QS, 12 play cups taken down,(not charged).  BAS and eggs,  room and stores ok.  3 Supers,  Top Full. next , threequarters full, bottom filling but less in than last week as being used during bad weather. Temper 7.
Hive#5   Inspected by  Richard W,    Brood chamber: Deep and a half,  QNS,  No eggs but a few sealed brood left, No Qc’s.   2 Qc’s from Hive 6, on frame with brood – no bees, put in. (nice calm bees).   Leave Till  10/7..    Temper 4.
Hive#6   Inspected by David P.    Brood chamber : Deep and a Half,   QS,  6 Qc’s,  BAS and new eggs.  stores and room OK.  Temper 9.    – moved 1 frame with 2 Qc’s on into Hive 5. Other Qc’s taken down.  Replaced with frame of foundation, giving more room in brood nest.
Hive#7  Inspected by Richard W.   Brood Chamber: Deep.   QNS,  3 open hatched Qc’s.  no brood or eggs. ( poss virgin Q ), stores room OK, 2 Supers, Top full.   Temper 9.  bees very calm so poss Virgin Q present.  Put back together ASAP.
Hive#7A  Inspected by Richard W.   Brood Chamber 3 Shallows.    QNS.  3 open Hatched Qc’s. but 4 shallow frames of 4 – 5 day old brood, and some eggs. so new Q now laying up.  Stores and room OK,  Bottom super MT, so put on top of QX.     Temper 9.
Hive#8    Not gone through as Qc’s left on 12/6.   – Check again  3/7.
Hive#10   Top Bar Hive.     Inspected by Liz/Richard,     QS,  No Qc’s,   4 f BAS. and eggs. 2 stores room OK.   Temper 10.  – Ants still a problem nesting between the top bars.   Must bring a hoover to clear hive debris from OMF.!!!
Hive# 15 F   Inspected by Liz/Richard.   Brood Chamber: Deep,  QNS, no Qc’s,  5 f  BAS and 1 f eggs,  3 stores 2 undrawn  f for space.  added super above QX for added space. Temper 10.
Hive#15 H  Inspected By Liz.   Brood Chamber: Deep.  QNS, few play cups.  5 f BAS, stores and room OK,  1 Super starting to fill.  Temper 7,    Moved 3 Feet towards space at no3, After Inspection, but bees still a bit “grotty”.( Liz’s term)
Nucs 16A and 16B  not gone through,  but area of ferns cleared from around Nuc’s.  Bees Flying.